Grade 12th passAlgebra

Simplify:
(35 × 105 × 25)/ (57 × 65 )
Simplify:
(35 × 105 × 25)/ (57 × 65 )

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

105 can be written as = (5 × 2)5
= 55 × 25
25 can be written as = 5 × 5
= 52
65 can be written as = (2 × 3)5
= 25 × 35
Then we have,
= (35 × 55 × 25 × 52)/ (57 × 25× 35)
= (35 × 55+2× 25)/ (57 × 25 × 35) ... [∵am × an = am + n]
= (35 × 57 × 25)/ (57 × 25 × 35)
= (35-5 × 57 – 7 × 25-5)
= (30 × 50 × 20) ... [∵am ÷ an = am – n]
= 1 × 1× 1
= 1

Thanks